# Physics-constrained convolutional neural networks for inverse problems in spatiotemporal partial differential equations

### v3

@@ 6.5 Robustness
First, in the linear convection-diffusion problem as shown in Figure 4(a), we demonstrate that the relative error is largely independent of the form of parameterised bias. The relative magnitude $\mathcal{M}$ and Rastrigin wavenumber $k_{\phi}$ have small impact on the ability of the model to uncover the solution to the partial differential equation, with the performance being consistent across all cases. The model performs best for the case $k_{\phi}=7,\mathcal{M}=0.5$, achieving a relative error of $e=2.568\times 10^{-1}$. Second, results for the nonlinear convection-diffusion case show marked improvement in comparison with the linear case, with errors for the numerical study shown in Figure 4(b). Despite the introduction of a nonlinear operator, the relative error is consistently lower regardless of the parameterisation of bias. The network remains equally capable of uncovering the true state across a wide range of modalities and magnitudes. Although the physics loss from the network predictions ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ no longer scales in a linear fashion, this is beneficial for the training dynamics. This is because the second order nonlinearity promotes convexity in the loss-surface, which is then exploited by our gradient-based optimisation approach. Third, the PC-CNN is able to uncover the true state of the turbulent flow. Results shown in Figure 4(c) demonstrate an improvement in performance from the standard nonlinear convection-diffusion case. For both fixed Rastrigin wavenumber and magnitude, increasing the value of the parameter tends to decrease the relative error. The relative error in Figure 4(c) decreases as the non-convexity of the bias increases. This is because it becomes increasingly simple to distinguish the bias from the underlying flow field if there is a larger discrepancy. We also emphasise that these tests were run with a fixed computational budget to ensure fair comparison – with additional training, we observe a further decrease in relative error across all experiments. (We also ran tests with biases parameterized with multiple wavenumbers. We found that our method robustly infers the biases — results not shown.)
@@ 7.1 Comparison with standard upsampling
We show the results for a scale factor $\kappa=7$, with results for further scale factors shown in C. Results are compared with interpolating upsampling methods, i.e. bi-linear, and bi-cubic interpolation to demonstrate the ability of the method. We quantify the accuracy by computing the relative error between the true solution, ${\bm{u}}({\bm{\Omega}_{H}},t)$, and the corresponding network realisation, ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$
@@
Upon discarding the transient, a solution ${\bm{u}}({\bm{\Omega}_{H}},t)\in\mathbb{R}^{70\times 70}$ is generated by time-integration over $12\times 10^{3}$ time-steps, with $\Delta t=1\times 10^{-3}$. We extract the low-resolution solution ${\bm{u}}({\bm{\Omega}_{L}},t)\in\mathbb{R}^{10\times 10}$. We extract $2048$ samples at random from the time-domain of the solution, each sample consisting of $\tau=2$ consecutive time-steps. The adam optimiser (Kingma & Ba, 2015) is employed for training with a learning rate of $3\times 10^{-4}$. We take $\lambda_{P}=10^{3}$ as the regularisation factor for the loss, as shown in Eq. (4.1), and train for a total of $10^{3}$ epochs, which is empirically determined to provide sufficient convergence, as per the results of a hyperparameter study (result not shown).
@@
Although the relative error provides a notion of predictive accuracy, it is crucial to assess the physical characteristics of the reconstructed field (Pope & Pope, 2000). The energy spectrum, which is characteristic of turbulent flows, represents a multi-scale phenomenon where energy content decreases with the wavenumber. From the energy spectrum of the network’s prediction, ${\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega}_{L},t))$, we gain physical insight into the multi-scale nature of the solution. Results in Figure 7 show that the energy content of the low-resolution field diverges from that of the high-resolution field, which is a consequence of spectral aliasing. Network realisations ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$ are capable of capturing finer scales of turbulence compared to both interpolation approaches, prior to diverging from the true spectrum as $\lvert{\bm{k}}\rvert=18$. These results show the efficacy of the method, recovering in excess of $99\%$ of the system’s total energy. The physics loss, $\mathcal{L}_{P}$, enables the network to act beyond simple interpolation. The network is capable of de-aliasing, thereby inferring unresolved physics. Parametric studies show similar results across a range of scale factors $\kappa$ (shown in C).
@@ A.2 Reconstruction from sparse information
For the reconstruction task, a three-layer convolutional neural network is used for the model. Figure 9 provides an overview of the architecture, as well as the general training loop. The baseline network configuration is similar to that of Dong et al. (2014), with the additional constraints imposed through the physics-based loss function. The input and output layers each feature two channels, corresponding to the two velocity components of the flow being studied.
@@ Appendix B Pseudo-spectral discretisation of Navier-Stokes residual
Operating in the Fourier domain eliminates the diverge-free constraint (Canuto et al., 1988). The equations for the spectral representation of the Kolmogorov flow are
@@
with ${\bm{\hat{f}}}_{k}=-\left(\widehat{{\bm{u}}\cdot\nabla{\bm{u}}}\right)_{k}$, where nonlinear terms are handled pseudospectrally, employing the $\nicefrac{{2}}{{3}}$ de-aliasing rule to avoid unphysical culmination of energy at the high frequencies Canuto et al. (1988). A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, choosing a time-step $\Delta t$ that satisfies the Courant-Friedrichs-Lewy (CFL) condition. Initial conditions are generated by producing a random field scaled by the wavenumber, which retains the spatial structures of varying lengthscale in the physical domain (Ruan & McLaughlin, 1998).
Appendix C Parametric study on the scale factor

### v4

@@ 6.5 Robustness
First, in the linear convection-diffusion problem as shown in Figure 4(a), we demonstrate that the relative error is largely independent of the form of parameterised bias. The relative magnitude $\mathcal{M}$ and Rastrigin wavenumber $k_{\phi}$ have small impact on the ability of the model to uncover the solution to the partial differential equation, with the performance being consistent across all cases. The model performs best for the case $k_{\phi}=7,\mathcal{M}=0.5$, achieving a relative error of $e=2.568\times 10^{-1}$. Second, results for the nonlinear convection-diffusion case show marked improvement in comparison with the linear case, with errors for the numerical study shown in Figure 4(b). Despite the introduction of a nonlinear operator, the relative error is consistently lower regardless of the parameterisation of bias. The network remains equally capable of uncovering the true state across a wide range of modalities and magnitudes. Although the physics loss from the network predictions ${\bm{\eta_{\theta}}}({\bm{\zeta}}(\bm{\Omega},t))$ no longer scales in a linear fashion, this is beneficial for the training dynamics. This is because the second order nonlinearity promotes convexity in the loss-surface, which is then exploited by our gradient-based optimisation approach. Third, the PC-CNN is able to uncover the true state of the turbulent flow. Results shown in Figure 4(c) demonstrate an improvement in performance from the standard nonlinear convection-diffusion case. For both fixed Rastrigin wavenumber and magnitude, increasing the value of the parameter tends to decrease the relative error. The relative error in Figure 4(c) decreases as the non-convexity of the bias increases. This is because it becomes increasingly simple to distinguish the bias from the underlying flow field if there is a larger discrepancy. We also emphasise that these tests were run with a fixed computational budget to ensure fair comparison – with additional training, we observe a further decrease in relative error across all experiments. (We also ran tests with biases parameterized with multiple wavenumbers, an example of which is shown in Appendix C).
@@ 7.1 Comparison with standard upsampling
We show the results for a scale factor $\kappa=7$, with results for further scale factors shown in D. Results are compared with interpolating upsampling methods, i.e. bi-linear, and bi-cubic interpolation to demonstrate the ability of the method. We quantify the accuracy by computing the relative error between the true solution, ${\bm{u}}({\bm{\Omega}_{H}},t)$, and the corresponding network realisation, ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$
@@
Upon discarding the transient, a solution ${\bm{u}}({\bm{\Omega}_{H}},t)\in\mathbb{R}^{70\times 70}$ is generated by time-integration over $12\times 10^{3}$ time-steps, with $\Delta t=1\times 10^{-3}$. We extract the low-resolution solution ${\bm{u}}({\bm{\Omega}_{L}},t)\in\mathbb{R}^{10\times 10}$. We extract $2048$ samples at random from the time-domain of the solution, each sample consisting of $\tau=2$ consecutive time-steps. The adam optimiser (Kingma and Ba,, 2015) is employed for training with a learning rate of $3\times 10^{-4}$. We take $\lambda_{P}=10^{3}$ as the regularisation factor for the loss, as shown in Eq. (4.1), and train for a total of $10^{3}$ epochs, which is empirically determined to provide sufficient convergence, as per the results of a hyperparameter study (result not shown).
@@
Although the relative error provides a notion of predictive accuracy, it is crucial to assess the physical characteristics of the reconstructed field (Pope and Pope,, 2000). The energy spectrum, which is characteristic of turbulent flows, represents a multi-scale phenomenon where energy content decreases with the wavenumber. From the energy spectrum of the network’s prediction, ${\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega}_{L},t))$, we gain physical insight into the multi-scale nature of the solution. Results in Figure 7 show that the energy content of the low-resolution field diverges from that of the high-resolution field, which is a consequence of spectral aliasing. Network realisations ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$ are capable of capturing finer scales of turbulence compared to both interpolation approaches, prior to diverging from the true spectrum as $\lvert{\bm{k}}\rvert=18$. These results show the efficacy of the method, recovering in excess of $99\%$ of the system’s total energy. The physics loss, $\mathcal{L}_{P}$, enables the network to act beyond simple interpolation. The network is capable of de-aliasing, thereby inferring unresolved physics. Parametric studies show similar results across a range of scale factors $\kappa$ (shown in D).
@@ A.2 Reconstruction from sparse information
For the reconstruction task, a three-layer convolutional neural network is used for the model. Figure 9 provides an overview of the architecture, as well as the general training loop. The baseline network configuration is similar to that of Dong et al., (2014), with the additional constraints imposed through the physics-based loss function. The input and output layers each feature two channels, corresponding to the two velocity components of the flow being studied.
@@ Appendix B Pseudo-spectral discretisation of Navier-Stokes residual
Operating in the Fourier domain eliminates the diverge-free constraint (Canuto et al.,, 1988). The equations for the spectral representation of the Kolmogorov flow are
@@
with ${\bm{\hat{f}}}_{k}=-\left(\widehat{{\bm{u}}\cdot\nabla{\bm{u}}}\right)_{k}$, where nonlinear terms are handled pseudospectrally, employing the $\nicefrac{{2}}{{3}}$ de-aliasing rule to avoid unphysical culmination of energy at the high frequencies Canuto et al., (1988). A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, choosing a time-step $\Delta t$ that satisfies the Courant-Friedrichs-Lewy (CFL) condition. Initial conditions are generated by producing a random field scaled by the wavenumber, which retains the spatial structures of varying lengthscale in the physical domain (Ruan and McLaughlin,, 1998).
Appendix C Multi-wave number bias removal
Further to the parameterisation outlined in Section 6.1, we consider stationary biases that contain multiple wave numbers. We define this composite bias as the mean of two bias fields, with $k_{\phi}\in\{3,7\},\mathcal{M}=0.5$. Bias removal was performed for the Kolmogorov flow case, following the same procedure as outlined in Section 2.1. Results for the bias removal achieved a relative error of $5.83\times 10^{-2}$, with results for multiple time-steps shown in Figure 10. The structure of the predicted bias captures the extrema of the bias, however there is small discrepancy in the magnitude as the field should be symmetric along both axes. This further demonstrates the robustness of the proposed approach.
[…]
Figure 10: Removal of a stationary bias field containing multiple wave numbers. The first three columns show the biased, original, and predicted flow fields respectively. The rightmost column depicts the predicted additive bias. Results are shown for multiple time-steps in the simulation trajectory.
Appendix D Parametric study on the scale factor